Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Friday, November 6, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 6, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/06/98 (Friday)

Segment 1. Report for 11/06/98: The 8 mile segment of pipe from EMP 33.34 to EMP 25.48 was filled with water again in prepartation for hydrotest. Station construction and electrical work continued at Emido and Grapevine Station sites. Pipe was delivered and strung near EMP .05. The welding and NDT crews worked along this section. Trench was backfilled from EMP 7.05 back along Aquaduct access road B, near EMP 6.4. A machine ground asphalt in front of the trenching operation on Grapevine Rd, where the excavator proceeded south near EMP 16. Water trucks worked along the ROW for compaction and dust suppression. The trenching operation at the tie-in site at Lebec and Fazier Mtn Rds (~EMP 25) was halted when a 2" waterline was exposed and began to sag. The utility company was notified and work was halted until the utility was stabilized with a galvanized sleeve. Pipe bending took place near EMP 18.19. An Excavator opened trench on the Tejon Bench near EMP 18.50. Tie-in work on Tejon benchtook place near EMP16.91. Work continued at the I-5 bore area EMP 14.62.

Segment 2 (ANF to Sylmar) Report for 11/6/98: 2C (Templin Hwy to Castaic): Construction of a block wall on the ANF access road south of Templin Hwy continued. Some clean-up and materials removal occurred along this access road. Revegetation continued at several locations between stream crossing #46, EMP 53.00, and the testing cutout at EMP 55.46. Welding reached EMP 55.48, and is now complete for Segment 2C except for two testing cutouts. Coating moved north to EMP 55.70, shading to 55.87, F/O conduit laying to 55.93, and backfilling to 56.04. 2B (Castaic to Sylmar): At Feedmill Road, set-up of the bore drill was disrupted by mechanical breakdown. Backfill continued at Calgrove on both bore pits. 300 feet of sawcutting of the roadway of The Old Road was accomplished to EMP 72.3. 250 feet of trench was excavated south of Sierra Hwy to EMP 72.76. Bending, lowering-in, welding, x-ray, and coating occurred near Sunshine Cyn. 300 feet of asphalt and 120 feet of slurry were completed to EMP 73.2. General clean-up occurred at Weldon Cyn.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 6 November 1998: 3A: NIGHT crews finished hanging the bridge crossing over San Fernando Road, MP 86.77. Crews are scheduled to work on Hollywood Way tonight, MP 85.60. Electrical work continued on the block valve north of the Burbank Animal Shelter, MP 87.42. Excavation continued for the second day on the 33 feet deep bore pit on the south side of Tuxford Street. Crews at this location, MP 83.21, also worked on preparing shields for the bore pit. Backfill crews continued cleaning up the right of way on either side of the Verdugo Wash Bridge crossing today, MP 91.09. Pipeline marker locations continued to be staked out across Segment 3. 3B: Crews completed installing razor wire above the Downey and Riverside Bridges, ~MP 97.20. All materials were removed from the Los Angeles River channel. Clean up crews began their work on the north end of Downey Bridge, at the edge of the MTA maintenance facility. 3C: The Hubbard Street bore kicked off today, with the crew achieving 60 feet of casing pipe, MP 77.33. The six inch bore for the FOC at Roxford Street was completed today, after the pipe breaking twice during pulling, MP 75.42. Backfill continued their clean up and grading of the right of way near Grapevine Creek, completing work up to MP 74.73 (with skips). Welding crews worked on the mainline near MP 74.33, at the Hubbard bore, as well as on the casing pipe at Roxford Street. X-ray and coating crews worked on the mainline to near MP 74.30. The crew at the California Aqueduct crossing, MP 73.90, pulled the forms down from their concrete pour yesterday.

Segment 4B (El Segundo to Lynwood) 11-6-98: Backfilling activities continued at 108 Th St. ESEMP 1.11, Santa Ana ESEMP 0.04, and backfilling at Alameda and Fernwood AV ESEMP 0.02 has been completed. Construction activities continued at El Segundo station ESEMP 11.00. This crew was pushing the pipe through the casing. The crew at Slater ESEMP 1.68 was excavating to repair the conduit for the fiberoptics. The crew at Lynwood station EMP 0.0. was getting ready for the hydro test inside the station. The crew at 109 Th St. ESEMP 1.23 was testing the fiberoptics conduit.

Segment 5 (Lynwood to Long Beach) Report for 6 November 1998: Horizontal directional drill is pulling reamer through without any problems. The bores at the Wilmington Station are not encountering any problems.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 5 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page